Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 14

1. Section 6.3, problems 25, 28. Section 6.4, problm 23.

Figure 1: Section 6.3, problems 25, 28. Section 6.4, problm 23